OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32 A. an = -4 + an-1 B. an = -2 + an-1 C. an = -2 • an-1 D. an = -4 • an-1

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (campbell_st):

well a term in a geometric series is \[a_{n} = a_{1} r^{n -1}\] you have the 1st term -2 you need to find the common ratio r... is can be done by comparing terms \[\frac{a_{2}}{a_{1}} = \frac{a_{3}}{a_{2}}\] hope this helps

OpenStudy (anonymous):

I am a parent trying to help my child and I don't understand this.... she has the answer but I don't know if its right

OpenStudy (campbell_st):

what answer does she have..?

OpenStudy (anonymous):

c

OpenStudy (campbell_st):

each term is being multiplied by -4 so term 2 = -4 x term 1 8 = -4 x -2 if its correct try it with term 3 term 3 = -4 x term 2 -32 = -4 x 8 is it correct.. hope this helps...

OpenStudy (anonymous):

so is C right?

OpenStudy (anonymous):

\[(-1)^{n} * (2)^{2n-1}\] is useful? i can solve 99 percent of math questions

hero (hero):

99% are you sure @mahdi2020?

OpenStudy (anonymous):

yes sure i`m a engeenier

OpenStudy (anonymous):

ok but I don't understand what is the right answer?

OpenStudy (anonymous):

@Hero can you please help me with this answer

OpenStudy (anonymous):

you can check the answers input n=1, 2 3 , ... my ans is ok